OpenStudy (anonymous):

Suppose a, b, and c ∈ Z and that gcd(a, b) = 1. If a | bc, then what is the gcd(a, c) and the gcd (b, c) ?

OpenStudy (rational):

a | bc and gcd(a, b) = 1 => a | c => gcd(a, c) = a

OpenStudy (rational):

for second part : a|c => c = ak gcd(b, c) = gcd(b, ak) = gcd(b, k) = gcd(b, c/a)

OpenStudy (anonymous):

a|c means a is a divisor of c ... correct? and a | bc means a is a divisor/factor of b * c? not even sure of notation? any suggestions of resources to better understand this notation?

OpenStudy (rational):

you're right ! you can also "read" it as : a | b means a divides b

OpenStudy (rational):

where ever you see "|", replace it with the word "divides"

OpenStudy (anonymous):

a divides b ? gawd...why has it taken me so long to see that notation?

OpenStudy (rational):

yes to be more accurate : a|b means : "a divides b" evenly without any remainder...

OpenStudy (anonymous):

does it mean "divides evenly" you read my mind:)

OpenStudy (rational):

or "a goes into b" evenly without any remainder

OpenStudy (anonymous):

how do I give you a medal:)

OpenStudy (rational):

yep ! but i read it as "a divides b" ... evenly is implicit :)

OpenStudy (anonymous):

thank you so much!

OpenStudy (rational):

refer to 2nd chapter

OpenStudy (anonymous):

just started Abstract Algebra...I have heard it is a very hard course...even harder than linear alg or calc 2

OpenStudy (rational):

oh i never took abstract algebra before.... im into engineering... but i dont think its tougher than analysis lol..

OpenStudy (anonymous):

some day...i will take analysis:)

OpenStudy (rational):

for me number theory itself is hard lol, let alone even think of abstract/analysis -.-

OpenStudy (rational):

most ppl who take NT, abstract and analysis claim NT is the easy one :/

OpenStudy (anonymous):

which of these must be true of my first question? it has to be b right? a. gcd(a, c) = 1. b. gcd(a, c) ≠ 1. c. gcd (b, c) = 1. d. gcd(b, c) ≠ 1.

OpenStudy (rational):

Yep !

OpenStudy (rational):

by any chance is it mentioned that \(a \ne 1\) for this question anywhere ?

OpenStudy (anonymous):

thanks again man:)

OpenStudy (anonymous):

no...it is not mention anywhere...that \[a \neq 1\]

OpenStudy (anonymous):

I think that is why they put MUST in the question...

OpenStudy (anonymous):

@rational did you ask me for help with something...I was at lunch